OpenStudy (anonymous):

12/square root of 15 Simplify, but keep in radical form.

OpenStudy (anonymous):

\[12\div \sqrt{15}\]

OpenStudy (ghazi):

12/3.872=3.099

OpenStudy (anonymous):

I have to keep it in radical form, so is it already simplified?

OpenStudy (ghazi):

okay... \[12= 2*\sqrt{3}.....\frac{ 2*\sqrt{3} }{ \sqrt{3}\sqrt{5} }\]

OpenStudy (ghazi):

can you do it now...?

OpenStudy (anonymous):

alright, I see it. Thanks :)

OpenStudy (ghazi):

:)

OpenStudy (asnaseer):

@ghazi how did you get: \(12=2*\sqrt{3}\)

OpenStudy (asnaseer):

this would be true only for: \(\sqrt{12}=\sqrt{4*3}=2\sqrt{3}\)

OpenStudy (ghazi):

yea...i see it...it seems i have made a mistake...exactly @anaseer you're right

OpenStudy (asnaseer):

@cocoa020 what you need to do here is multiply the numerator and the denominator by \(\sqrt{15}\) and then simplify

OpenStudy (ghazi):

i didn't see it's not under square root

OpenStudy (asnaseer):

so:\[\frac{12}{\sqrt{15}}=\frac{12}{\sqrt{15}}\times\frac{\sqrt{15}}{\sqrt{15}}=\frac{12\sqrt{15}}{15}\]

OpenStudy (asnaseer):

the idea is to ensure you don't have any radicals in the denominator

OpenStudy (asnaseer):

does that make sense @cocoa020 ?